Greta Cook died in 2018. Her gross estate was $11,290,000, administration expenses of $60,000 and an adjusted taxable gifts were $20,000. Compute the estate tax liability.

I am able to find the estate tax base of $11,250,000, but I am unsure of how to find the tax liability.
Answer #1

Given,

Gross Estate = $11290000

Administration expense = $60000

Adjusted taxable gifts = $20000

Net Estate = Gross Estate - Administration Expense

= 11290000 - 60000

= $11230000

For 2018, Estate tax exemption is $11180000. But the value of taxable gifts is subtracted to find out the final exemption.

Net exemption = 11180000 - 20000

= $11160000

Taxable Estate = Net estate - Net exemption

= 11230000 - 11160000

= $70000

For 2018, Estate Tax rate is 40%. Therefore,

Estate tax liability = 40% of $70000

= 0.4 X 70000

= $28000

Note: I have subtracted 20000 worth gifts from exemption. In stead if you add it in the net estate, still the result will be same. But as per the rule, the method as shown above is followed.